### v1

# ExteriorExtensions: A package for Macaulay2

Abstract.
We explain a Macaulay2 implementation of a construction, which appeared in [HolweckOeding22], of a graded algebra structure on the direct sum of a Lie algebra $\mathfrak{g}$ (typically $\mathfrak{sl}_{n}$) and a $\mathfrak{g}$-module (typically a subspace of an exterior algebra ${\textstyle\bigwedge^{\bullet}}\mathbb{C}^{n}$). We implement brackets, a Killing form, matrix representations of adjoint operators and ranks of blocks of their powers.
1. Introduction
The Lie algebra $\mathfrak{g}=\mathfrak{sl}_{n}(\mathbb{F})$ is a matrix algebra with bracket $[A,B]=AB-BA$ for traceless matrices $A,B\in\mathbb{F}^{n\times n}$. The package ExteriorExtensions extends this algebra to a larger one $\mathfrak{a}=\mathfrak{g}\oplus M$ for a $\mathfrak{g}$-module $M$, and at this time we have only implemented this for $M$ a sum of exterior powers of the underlying vector space $V=\mathbb{F}^{n}$. We say that the extension algebra has GJD (Jordan decomposition consistent with the $G$-action) if the $G=\text{Lie}(\mathfrak{g})$-action on $M$ is consistent with conjugation on adjoint operators $\operatorname{ad}_{T}\colon\mathfrak{a}\to\mathfrak{a}$ (see [HolweckOeding22]). The package provides functionality for the standard matrix representations of adjoint operators on this algebra, as well as a tabulation of the ranks of blocks of powers of adjoint operators. We also compute the Killing matrix, which may help in identifying the algebra.
1.1. Exterior algebras
The exterior algebra on a vector space $V$ is the direct sum of exterior powers ${\textstyle\bigwedge^{\bullet}}V=\bigoplus_{k\geq 0}{\textstyle\bigwedge^{k}}V$, with product given by concatenation, which respects a $\mathbb{Z}$-grading. The exterior algebra on $V$ is Artinian when $V$ is finite dimensional. However, we can obtain a $\mathbb{Z}_{n}$-grading if $n=\dim V$ by choosing a volume form $0\neq\Omega\in{\textstyle\bigwedge^{n}}V$ and replacing products $\varphi\wedge\psi$ with $\Omega(\Omega(\varphi)\wedge\Omega(\psi))$ if $\deg(\varphi)+\deg(\psi)>m$. Moreover, we note that one can obtain a (possibly) different grading other than the degree grading by declaring ${\textstyle\bigwedge^{j}}V$ to be the grade-1 piece, and defining the rest of the graded pieces inductively.
1.2. Extending the Lie algebra by appending an exterior algebra
Now we consider the graded vector space $\mathfrak{a}=\mathfrak{g}\oplus{\textstyle\bigwedge^{k}}V\oplus{\textstyle\bigwedge^{2k}}V\oplus\dots=\mathfrak{g}\oplus({\textstyle\bigwedge^{\bullet}}V\setminus{\textstyle\bigwedge^{0}}V)$, where we delete the scalars from the exterior algebra, and set $\mathfrak{a}_{0}=\mathfrak{g}$ and $\mathfrak{a}_{i}={\textstyle\bigwedge^{ik\mod n}}V$ for $1\leq i<n=\dim V$.
Recall that a structure tensor $B\in\mathfrak{a}^{*}\otimes\mathfrak{a}^{*}\otimes\mathfrak{a}$ representing the bilinear product $[\;,\;]\colon\mathfrak{a}\times\mathfrak{a}\to\mathfrak{a}$ is equivalent information to the algebra structure on $\mathfrak{a}$. To respect the grading we must require that $B_{i,j,k}=0$ whenever $i+j\not\equiv k\mod n$. In [HolweckOeding22, Theorem 3.12] we showed that requiring GJD implies that there is an essentially unique structure tensor (product) on $\mathfrak{a}$. We implement that product (a method function in M2 called bracket) as follows.
1.2.1. The structure tensor / bracket
For $B_{0,0,0}$ we have the standard skew-commuting bracket on $\mathfrak{sl}_{n}$, namely $[A,B]=AB-BA$. For $B_{0,j,j}$ (and for $B_{j,0,j}$) with $j>0$ we use the standard Lie algebra action of $\mathfrak{sl}_{n}$ on the fundamental representation ${\textstyle\bigwedge^{k_{j}}}\mathbb{C}^{n}$, namely for $A\in\mathfrak{sl}_{n}$ and $T\in{\textstyle\bigwedge^{k}}V$ we take $[A,T]:=A.T$, where . denotes the Lie algebra action. Similarly $[T,A]=-A.T$.
[…]
$\begin{matrix}\mathfrak{a}_{i}\times\mathfrak{a}_{j}&\to&\mathfrak{a}_{0}\\ (S,T)&\mapsto&[S,T]=-k\pi\left(\sum_{p,q}\left(\frac{\partial S}{\partial e_{p}}\lrcorner\frac{\partial*T}{\partial e_{q}}\right)\right)\end{matrix}\quad\quad\text{when }i+j=n\text{ and }j>n/2.$
[…]
1.2.2. Adjoint Operators
The standard matrix representation of the adjoint operator on $\mathfrak{a}$ defined by $\operatorname{Ad}_{x}(y)=[x,y]$ is constructed by choosing an ordered basis of $\mathfrak{a}$, and expressing the results of $\operatorname{Ad}_{x}(y)$ for each basis vector $y$ in terms of the standard basis as the corresponding column. One caveat is that we have to choose a basis of $\mathfrak{sl}_{n}$, and there is not a "standard" choice for the basis of the Cartan subalgebra of traceless diagonal matrices. We choose a common one $\{E_{ii}-E_{i+1,i+1}\mid i=1\ldots n-1\}$. Then we have to solve a system of equations to find the coordinates for the cartan rather than just extracting the entries of the matrix.
Respecting the grading yields block structure given by $\operatorname{Ad}_{x}(y)=(B_{i,j,i+j\mod n}(y))_{i,j}$. The ranks of the blocks of $\operatorname{Ad}_{x}$ are invariants of $x$, as are the ranks of blocks of the powers of $\operatorname{Ad}_{x}$. We collect these ranks in a table and display them with a function prettyBlockRanks.
The Jacobi identity is stated via adjoint operators as $[\operatorname{Ad}_{x},\operatorname{Ad}_{y}]=\operatorname{Ad}_{[x,y]}$ [FultonHarris].
1.2.3. The Killing form
[…]
2. Examples
The algebras $\mathfrak{a}=\mathfrak{sl}_{2m}\oplus{\textstyle\bigwedge^{m}}\mathbb{C}^{2m}$ for $m=2$ and $m=4$ are actually simple Lie algebras, respectively $\mathfrak{sl}_{4}\oplus{\textstyle\bigwedge^{2}}\mathbb{C}^{4}\cong\mathfrak{sp}(6)$ and $\mathfrak{sl}_{8}\oplus{\textstyle\bigwedge^{4}}\mathbb{C}^{8}\cong\mathfrak{e}_{7}$. An excellent overview on fine gradings of simple Lie algebras is given in [MR3601079]. It is not immediately clear to the author if this example falls within those fine gradings classified in [MR3601079, Example 6].
Example 1.
We have a sporadic isomorphism $\mathfrak{sl}_{4}\oplus{\textstyle\bigwedge^{2}}\mathbb{C}^{4}\cong\mathfrak{sp}_{6}$. Recall from [FultonHarris, Ch. 15.2 and Ch.17.1] the structure of the two algebras. The weights of ${\textstyle\bigwedge^{2}}\mathbb{C}^{4}$ are the vertices of an octahedron, while the weights of $\mathfrak{sl}_{4}$ are the midpoints of the edges of a cube. These geometries are also present in the weights of a basis of $\mathfrak{sp}_{6}$, see [FultonHarris, p220, p254]. We set the algebra as an ExteriorExtension via:
i1 : loadPackage"ExteriorExtensions"
o1 = ExteriorExtensions
o1 : Package
i2 : ea24 = buildAlgebra(2,4,e,QQ)
o2 = ea24
o2 : ExteriorExtension
We will see that $\mathfrak{sl}_{4}\oplus{\textstyle\bigwedge^{2}}\mathbb{C}^{4}$ with the (essentially unique) bracket coming from the GJD criterion forms a Lie algebra. We can check (at least on random points) if the bracket satisfies the Lie algebra axioms in each graded piece of the algebra. For grade 0 times grade 0 we perform the following:
i3 : A = makeTraceless random(QQ^4,QQ^4)
o3 = | 443/168 9/2    3/2      3/5    |
     | 4       23/168 4        3/7    |
     | 3/2     5/2    -265/168 9/5    |
     | 9       1/3    5/8      -67/56 |
              4        4
o3 : Matrix QQ  <--- QQ
i4 : B = makeTraceless random(QQ^4,QQ^4)
o4 = | 13/280 1/4     3/5     2/3     |
     | 2/3    -71/280 1/10    3/10    |
     | 1/5    5/2     -51/280 5/3     |
     | 1      2/5     3/5     109/280 |
              4        4
o4 : Matrix QQ  <--- QQ
i5 : ea24.bracket(A,B) + ea24.bracket(B,A)
o5 = 0
              4        4
o5 : Matrix QQ  <--- QQ
i6 : ea24.ad(ea24.bracket(A,B)) - ea24.bracket(ea24.ad(A),ea24.ad(B))
o6 = 0
              21        21
o6 : Matrix QQ   <--- QQ
We shouldn’t be surprised that we get 0 for both since the grade 0 piece is a Lie algebra.
For a pair of grade 1 elements we perform the following:
i7 : A = random(2, ea24.appendage)
     4       5       7               5
o7 = -e e  + -e e  + -e e  + 5e e  + -e e  + 2e e
     7 0 1   9 0 2   9 1 2     0 3   6 1 3     2 3
o7 : QQ[e ..e ]
         0   3
i8 : B = random(2, ea24.appendage)
     1       3       2       2       1       10
o8 = -e e  + -e e  + -e e  + -e e  + -e e  + --e e
     5 0 1   2 0 2   3 1 2   9 0 3   3 1 3    3 2 3
o8 : QQ[e ..e ]
         0   3
i9 : ea24.bracket(A,B) + ea24.bracket(B,A)
o9 = 0
                        4                  4
o9 : Matrix (QQ[e ..e ])  <--- (QQ[e ..e ])
                 0   3              0   3
i10 : ea24.ad(ea24.bracket(A,B)) - ea24.bracket(ea24.ad(A),ea24.ad(B))
o10 = 0
               21        21
o10 : Matrix QQ   <--- QQ
For a point in grade 0 and a point in grade 1 we see the axioms hold, but we recall the caveat that it is important to make the matrix traceless, otherwise the Jacobi identity fails.
i11 : A = makeTraceless random(QQ^4,QQ^4)
o11 = | 17/80 10/9  7/2    1/2  |
      | 5/9   21/80 3/7    9/5  |
      | 1     4/5   -43/80 3/2  |
      | 1/4   3/2   3      1/16 |
               4        4
o11 : Matrix QQ  <--- QQ
i12 : B = random(2, ea24.appendage)
             9       2              8       10
o12 = e e  + -e e  + -e e  + e e  + -e e  + --e e
       0 1   5 0 2   9 1 2    0 3   9 1 3    9 2 3
o12 : QQ[e ..e ]
          0   3
i13 : ea24.bracket(A,B) + ea24.bracket(B,A)
o13 = 0
o13 : QQ[e ..e ]
          0   3
i14 : ea24.ad(ea24.bracket(A,B)) - ea24.bracket(ea24.ad(A),ea24.ad(B))
o14 = 0
               21        21
o14 : Matrix QQ   <--- QQ
We can also construct a Killing matrix and see that it is $21\times 21$ and non-degenerate. The only possibility for a simple Lie algebra over $\mathbb{C}$ of dimension 21 is $\mathfrak{sp}_{6}$. One checks that the Killing matrix has a full set of real eigenvectors, so the isomorphism $\mathfrak{a}\cong\mathfrak{sp}_{6}$ holds over $\mathbb{R}$.
i15 : K = ea24.KillingMatrix()
o15 = | 20 10 0  0  0  0  0  0  0  0  0  0  0  0  0  0  0   0  0  0   0  |
      | 10 20 10 0  0  0  0  0  0  0  0  0  0  0  0  0  0   0  0  0   0  |
      | 0  10 20 0  0  0  0  0  0  0  0  0  0  0  0  0  0   0  0  0   0  |
      | 0  0  0  0  0  0  0  0  0  10 0  0  0  0  0  0  0   0  0  0   0  |
      | 0  0  0  0  0  0  0  0  0  0  10 0  0  0  0  0  0   0  0  0   0  |
      | 0  0  0  0  0  0  0  0  0  0  0  10 0  0  0  0  0   0  0  0   0  |
      | 0  0  0  0  0  0  0  0  0  0  0  0  10 0  0  0  0   0  0  0   0  |
      | 0  0  0  0  0  0  0  0  0  0  0  0  0  10 0  0  0   0  0  0   0  |
      | 0  0  0  0  0  0  0  0  0  0  0  0  0  0  10 0  0   0  0  0   0  |
      | 0  0  0  10 0  0  0  0  0  0  0  0  0  0  0  0  0   0  0  0   0  |
      | 0  0  0  0  10 0  0  0  0  0  0  0  0  0  0  0  0   0  0  0   0  |
      | 0  0  0  0  0  10 0  0  0  0  0  0  0  0  0  0  0   0  0  0   0  |
      | 0  0  0  0  0  0  10 0  0  0  0  0  0  0  0  0  0   0  0  0   0  |
      | 0  0  0  0  0  0  0  10 0  0  0  0  0  0  0  0  0   0  0  0   0  |
      | 0  0  0  0  0  0  0  0  10 0  0  0  0  0  0  0  0   0  0  0   0  |
      | 0  0  0  0  0  0  0  0  0  0  0  0  0  0  0  0  0   0  0  0   20 |
      | 0  0  0  0  0  0  0  0  0  0  0  0  0  0  0  0  0   0  0  -20 0  |
      | 0  0  0  0  0  0  0  0  0  0  0  0  0  0  0  0  0   0  20 0   0  |
      | 0  0  0  0  0  0  0  0  0  0  0  0  0  0  0  0  0   20 0  0   0  |
      | 0  0  0  0  0  0  0  0  0  0  0  0  0  0  0  0  -20 0  0  0   0  |
      | 0  0  0  0  0  0  0  0  0  0  0  0  0  0  0  20 0   0  0  0   0  |
               21        21
o15 : Matrix QQ   <--- QQ
i16 : rank K
o16 = 21
i17 :  toList eigenvalues K
o17 =  {34.1421, 20, 5.85786, 10, -10, 10, -10, 10, -10, 10,
          -10, 10, -10, 10, -10, 20, -20, 20, -20, 20, -20}
o17 : List
The tensors $e_{i}\wedge e_{j}+e_{k}\wedge e_{l}$ for $i,j,k,l$ distinct have diagonalizable adjoint forms (we call them semi-simple or ad-semi-simple). We can check that the geometric and algebraic multiplicities of the eigenvalues add up:
i18 : A = ea24.ad(e_0*e_1 + e_2*e_3)
o18 = | 0 0  0 0 0  0 0  0 0 0 0 0 0  0  0 -1 0  0  0 0  1  |
      | 0 0  0 0 0  0 0  0 0 0 0 0 0  0  0 2  0  0  0 0  -2 |
      | 0 0  0 0 0  0 0  0 0 0 0 0 0  0  0 -1 0  0  0 0  1  |
      | 0 0  0 0 0  0 0  0 0 0 0 0 0  0  0 0  0  0  0 0  0  |
      | 0 0  0 0 0  0 0  0 0 0 0 0 0  0  0 0  0  -2 0 0  0  |
      | 0 0  0 0 0  0 0  0 0 0 0 0 0  0  0 0  2  0  0 0  0  |
      | 0 0  0 0 0  0 0  0 0 0 0 0 0  0  0 0  0  0  0 -2 0  |
      | 0 0  0 0 0  0 0  0 0 0 0 0 0  0  0 0  0  0  2 0  0  |
      | 0 0  0 0 0  0 0  0 0 0 0 0 0  0  0 0  0  0  0 0  0  |
      | 0 0  0 0 0  0 0  0 0 0 0 0 0  0  0 0  0  0  0 0  0  |
      | 0 0  0 0 0  0 0  0 0 0 0 0 0  0  0 0  0  0  2 0  0  |
      | 0 0  0 0 0  0 0  0 0 0 0 0 0  0  0 0  0  0  0 2  0  |
      | 0 0  0 0 0  0 0  0 0 0 0 0 0  0  0 0  -2 0  0 0  0  |
      | 0 0  0 0 0  0 0  0 0 0 0 0 0  0  0 0  0  -2 0 0  0  |
      | 0 0  0 0 0  0 0  0 0 0 0 0 0  0  0 0  0  0  0 0  0  |
      | 0 1  0 0 0  0 0  0 0 0 0 0 0  0  0 0  0  0  0 0  0  |
      | 0 0  0 0 0  1 0  0 0 0 0 0 -1 0  0 0  0  0  0 0  0  |
      | 0 0  0 0 -1 0 0  0 0 0 0 0 0  -1 0 0  0  0  0 0  0  |
      | 0 0  0 0 0  0 0  1 0 0 1 0 0  0  0 0  0  0  0 0  0  |
      | 0 0  0 0 0  0 -1 0 0 0 0 1 0  0  0 0  0  0  0 0  0  |
      | 0 -1 0 0 0  0 0  0 0 0 0 0 0  0  0 0  0  0  0 0  0  |
              21        21
o18 : Matrix QQ   <--- QQ
i19 : round\ toList eigenvalues A
o19 = {0, 0, 2, -2, -2, 0, 2, -2, 0, -2, 2, 2, -2, 2, 0, 0, 0, 0, 0, 0, 0}
o19 : List
i20 : rank A
o20 = 10
The semi-simple structure is also indicated by the block ranks computation:
i21 :  ea24.prettyBlockRanks A
o21 = |g00|g01|g10|g11|total|
      |0  |5  |5  |0  |10   |
      |5  |0  |0  |5  |10   |
      |0  |5  |5  |0  |10   |
      |5  |0  |0  |5  |10   |
The rows in the above table correspond to subsequent powers of the matrix.
We can check that neither the elements nor their adjoint forms commute, so they’re not simultaneously diagonalizable.
i22 : ea24.bracket(e_0*e_1 + e_2*e_3,e_0*e_2 + e_1*e_3)
o22 = {1} | 0 0  0  2 |
      {1} | 0 0  -2 0 |
      {1} | 0 -2 0  0 |
      {1} | 2 0  0  0 |
                         4                  4
o22 : Matrix (QQ[e ..e ])  <--- (QQ[e ..e ])
                  0   3              0   3
i23 : B = ea24.ad(e_0*e_2 + e_1*e_3);
               21        21
o23 : Matrix QQ   <--- QQ
i24 : 0*ea24.bracket(A,B)== ea24.bracket(A,B)
o24 = false
In fact, we checked that the only elements of ${\textstyle\bigwedge^{2}}\mathbb{C}^{4}$ that commute with $e_{i}\wedge e_{j}+e_{k}\wedge e_{l}$ are scalar multiples of itself. So we can’t hope for a Cartan subalgebra that intersects ${\textstyle\bigwedge^{2}}\mathbb{C}^{4}$ and includes more than one of these forms.
i25 : L = apply(ea24.bases#1, xx-> sub(ea24.bracket(xx,e_0*e_1 + e_2*e_3), QQ))
o25 = {| 1 0 0  0  |, | 0 0 0 -2 |, | 0 0 2 0 |, | 0  0 0 0  |, | 0  0 0 0 |, | -1 0  0 0 |}
       | 0 1 0  0  |  | 0 0 0 0  |  | 0 0 0 0 |  | 0  0 0 -2 |  | 0  0 2 0 |  | 0  -1 0 0 |
       | 0 0 -1 0  |  | 0 2 0 0  |  | 0 0 0 0 |  | -2 0 0 0  |  | 0  0 0 0 |  | 0  0  1 0 |
       | 0 0 0  -1 |  | 0 0 0 0  |  | 0 2 0 0 |  | 0  0 0 0  |  | -2 0 0 0 |  | 0  0  0 1 |
o25 : List
i26 : R = QQ[x_0.. x_5]
o26 = R
o26 : PolynomialRing
i27 : decompose ideal flatten entries sum(6, i-> sub(L_i, R)*x_i)
o27 = {ideal (x , x , x , x , x  - x )}
               4   3   2   1   0    5
o27 : List
i28 : first ea24.bases#1 + last ea24.bases#1
o28 = e e  + e e
       0 1    2 3
o28 : QQ[e ..e ]
          0   3
Example 2.
We consider $\mathfrak{sl}_{6}\oplus{\textstyle\bigwedge^{3}}\mathbb{C}^{6}$, which is not a Lie algebra as it fails to be skew-commuting and it fails the Jacobi identity.
i2 : ea36 = buildAlgebra(3,6,QQ);
i3 : A = random(3, ea36.appendage);
i4 : B = random(3, ea36.appendage);
i5 : ea36.bracket(A,B) + ea36.bracket(B,A) == 0
o5 = false
i6 : ea36.bracket(A,B) - ea36.bracket(B,A) == 0
o6 = true
i7 : ea36.ad(ea36.bracket(A,B)) - ea36.bracket(ea36.ad(A),ea36.ad(B)) == 0
o7 = false
Example 3.
Consider the case of $\mathfrak{sl}_{8}\oplus{\textstyle\bigwedge^{4}}\mathbb{C}^{8}\cong\mathfrak{e}_{7}$ considered in [Antonyan, oeding2022]. Antonyan shows that a 7-dimensional Cartan subalgebra is given by 4-forms of type $e_{I}+e_{J}$ with the concatenation $I|J$ a permutation of $[8]$, the 7 pairs of indices $I,J$ as follows:
$\begin{matrix}\{\{1,2,3,4\},\;\{5,6,7,8\}\},&&\{\{1,3,5,7\},\;\{6,8,2,4\}\},\\ \{\{1,5,6,2\},\;\{8,4,7,3\}\},&&\{\{1,6,8,3\},\;\{4,7,5,2\}\},\\ \{\{1,8,4,5\},\;\{7,2,6,3\}\},&&\{\{1,4,7,6\},\;\{2,3,8,5\}\},\\ \{\{1,7,2,8\},\;\{3,5,4,6\}\}\end{matrix}$
We can check the diagonalizability and commutativity as follows.
i2 : ea48 = buildAlgebra(4,8,e,QQ);
i3 : list2E = L-> sum(L, l-> product(l, i-> e_(i-1)));
i4 : aa = {{{1,2,3,4},{5,6,7,8}}, {{1,3,5,7},{6,8,2,4}},{{1,5,6,2},{8,4,7,3}},
     {{1,6,8,3},{4,7,5,2}},{{1,8,4,5},{7,2,6,3}},{{1,4,7,6},{2,3,8,5}},
     {{1,7,2,8},{3,5,4,6}}};
i5 : for i to length(aa)-1 do(
     adList_i = ea48.ad(list2E aa_i);
     print(# for xx in eigenvalues adList_i list if
     ¯not round(abs(xx)) ==0 then xx else continue, rank adList_i);     )
(66, 66)
(66, 66)
(66, 66)
(66, 66)
(66, 66)
(66, 66)
(66, 66)
i6 : matrix apply(7, i-> apply(7, j-> rank ea48.bracket(adList_i,adList_j) ))
o6 = 0
              7        7
o6 : Matrix ZZ  <--- ZZ
Example 4.
Larger algebras are also possible, such as $\mathfrak{e}_{8}\cong\mathfrak{sl}_{9}\oplus{\textstyle\bigwedge^{3}}\mathbb{C}^{9}\oplus{\textstyle\bigwedge^{6}}\mathbb{C}^{9}$ from [Vinberg-Elashvili], and we have done extensive computations in [HolweckOeding22]. Here are some example computations that check that the bracket satisfies the Lie-bracket conditions (skew-symmetry and Jacobi).
i2 :     ea = buildAlgebra(3,9)
o2 = ea
o2 : ExteriorExtension
i3 :     E = ea.appendage
o3 = E
o3 : PolynomialRing, 9 skew commutative variables
i4 :     A = makeTraceless random(QQ^9,QQ^9);
              9        9
o4 : Matrix QQ  <--- QQ
i5 :     B = makeTraceless random(QQ^9,QQ^9);
              9        9
o5 : Matrix QQ  <--- QQ
i6 :     ea.bracket(A,B) + ea.bracket(B,A)
o6 = 0
              9        9
o6 : Matrix QQ  <--- QQ
i7 :     ea.ad(ea.bracket(A,B)) -ea.bracket(ea.ad(A),ea.ad(B))
   A = makeTraceless random(QQ^9,QQ^9);
o7 = 0
              248        248
o7 : Matrix QQ    <--- QQ
i8 :
              9        9
o8 : Matrix QQ  <--- QQ
i9 :    B = random(3, E);
i10 :    ea.bracket(A,B) + ea.bracket(B,A)
o10 = 0
o10 : E
i11 :    t1 =    ea.ad(ea.bracket(A,B));
               248        248
o11 : Matrix QQ    <--- QQ
i12 :    t2 =  ea.bracket(ea.ad(A),ea.ad(B));
               248        248
o12 : Matrix QQ    <--- QQ
i13 :    t1-t2   -- gives zero!
o13 = 0
               248        248
o13 : Matrix QQ    <--- QQ
i14 :    A = makeTraceless random(QQ^9,QQ^9);
               9        9
o14 : Matrix QQ  <--- QQ
i15 :    B = random(6, E);
i16 :    ea.bracket(A,B) + ea.bracket(B,A)
o16 = 0
o16 : E
i17 :    ea.ad(ea.bracket(A,B)) - ea.bracket(ea.ad(A),ea.ad(B)) -- gives zero!
o17 = 0
               248        248
o17 : Matrix QQ    <--- QQ
i18 :    A = random(3, E);
i19 :    B = random(6, E);
i20 :    ea.bracket(A,B) + ea.bracket(B,A)
o20 = 0
              9       9
o20 : Matrix E  <--- E
i21 :    t1 =    ea.ad(ea.bracket(A,B));
               248        248
o21 : Matrix QQ    <--- QQ
i22 :    t2 =  ea.bracket(ea.ad(A),ea.ad(B));
               248        248
o22 : Matrix QQ    <--- QQ
i23 :    rank(t1-t2) -- gives zero!
o23 = 0
i24 :    A = makeTraceless random(QQ^9,QQ^9);
               9        9
o24 : Matrix QQ  <--- QQ
i25 :    pt = random(3, E);
During the development of this code we found a problem with our definition of the bracket that stemmed back to our implementation of the Hodge star operator, which we were able to fix. The following equations 2.5 and 2.6 from [Vinberg-Elashvili]
i26 :    ea.bracket(A, ea.star(pt )) + ea.star ea.bracket(transpose A, pt)
o26 = 0
o26 : E
i27 :    pt = E_0*E_1*E_2; tmp  = ea.bracket(pt, ea.star(pt))
o28 = {2} | -2/3 0    0    0   0   0   0   0   0   |
      {2} | 0    -2/3 0    0   0   0   0   0   0   |
      {2} | 0    0    -2/3 0   0   0   0   0   0   |
      {2} | 0    0    0    1/3 0   0   0   0   0   |
      {2} | 0    0    0    0   1/3 0   0   0   0   |
      {2} | 0    0    0    0   0   1/3 0   0   0   |
      {2} | 0    0    0    0   0   0   1/3 0   0   |
      {2} | 0    0    0    0   0   0   0   1/3 0   |
      {2} | 0    0    0    0   0   0   0   0   1/3 |
              9       9
o28 : Matrix E  <--- E
i29 :    ea.bracket(tmp, pt)
o29 = -2e e e
         0 1 2
o29 : E